OpenStudy (loser66):

calculate 1.80m + 142.5 cm+ 5.34* 10^5 \(\mu\)m=? Please help

OpenStudy (phi):

you should change to a common unit. Try meters

OpenStudy (loser66):

I tried \(\mu\)m

OpenStudy (psymon):

Well, we can just convert everything into meters and then add them. The mu m is micrometers and that is 1 millionth of a meter. So if you have 5.34x10^5, you have 534000 micrometers. But since micrometer is 1 millionth of a meter, you can say you have .534 meters. Once you have that meter conversion the others aren't that bad.

OpenStudy (loser66):

what is the answer?

OpenStudy (phi):

to change units, set up ratios for example, to change 142.5 cm to meters you know you want to multiply by a ratio with meters up top and cm in the bottom (so the cm will "cancel") how many cm in 1 meter ? 100 cm/1 m but we want that upside down, 1 m /100 cm now do 142.5 cm * 1 m /100 cm = 1.425 meters

OpenStudy (loser66):

I know how to convert. However, I am dealing with significant figures also. That's why I choose \(\mu\) to do.

OpenStudy (loser66):

here what I did 1.80 *10^6 +14.25*10^6 + 0.534*10^6= 1.65*10^7

OpenStudy (phi):

you should know there are 10e6 μm/ 1 m so 5.34* 10^5 μm * 1 m/10^6 μm= 5.34/10 m = 0.534 meters your problem is now 1.80m + 1.425 m + 0.534 m

OpenStudy (loser66):

You see, no matter what the unit we convert, the results should be the same. Ours are not.:(

OpenStudy (phi):

yes, and do you want to bet who is correct ? for starters, your conversion from cm to μm is incorrect.

OpenStudy (phi):

1.80m + 1.425 m + 0.534 m= 3.759 m

OpenStudy (loser66):

are you kidding me about "bet' ? If I have confidence on myself, I don't post that kind of question here, XD

OpenStudy (loser66):

however, what's wrong with my cm to \(\mu\)m?

OpenStudy (phi):

your cm->μm is 10x too big

OpenStudy (phi):

To me, it makes sense to convert to meters, because all the other distances are some factors of 10 different.

OpenStudy (loser66):

oh yes, I got my mistake, thanks a lot @phi

OpenStudy (loser66):

surely we should go the easiest way. I just want to practice the significant figure also.

OpenStudy (loser66):

to your result, for exactness, it should be 3.76m, right?

OpenStudy (phi):

In this case 1.80 m is the most uncertain measurement so we should round to the nearest 100th of a meter (cm) 3.76 m sounds good.

OpenStudy (loser66):

Thank you very much.